OpenStudy (anonymous):

how do you factor : a^2-2ab+b^2

OpenStudy (anonymous):

what times what = a^2? and what times what = b^2?

OpenStudy (anonymous):

axa=a^2 bxb=b^2 but what about the -2ab how do i get thet?

OpenStudy (anonymous):

when you factor a polynomial, and that's what that is, you get something like: (a + b) (c + d) if you are lucky and this expands to ac+ ad +bc + bd

OpenStudy (anonymous):

so what is (a - b) (a - b) ?

OpenStudy (anonymous):

ok now i got it I forgot that when you add -ab-ab= -2ab... because their two of them . thanks !

OpenStudy (anonymous):

that's all there is to it